OpenStudy (anonymous):

simplify by removing factors of 1: 6y-18 / 3-y :)

OpenStudy (anonymous):

go get 'em

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

(6y-18)/(3-y) = 6(y-3)/(3-y) = -6 (3-y)/(3-y)